OpenStudy (anonymous):

find all x such that: exp(x)=1

OpenStudy (anonymous):

exp() what does that mean?

OpenStudy (anonymous):

e^x = 1. Take the natural log on both sides: ln(e^x) = ln(1) x lne = 0 x = 0 :-)

OpenStudy (anonymous):

never really seen anyone write it like that.